OpenStudy (anonymous):

pleaseee simplify (2x+5)(3x-7)

OpenStudy (anonymous):

use F.O.I.L

OpenStudy (anonymous):

6 x^2+x-35

OpenStudy (anonymous):

(2x + 5) * (3x - 7) = (2x * 3x) + (2x * -7) + (5 * 3x) + (5 * -7) = 6x^2 - 14x + 15x - 35 = 6x^2 +x - 35